OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (texaschic101):

y - y1 = m(x - x1) <-- point slope form slope(m) = -3 (-1,6) x1 = -1 and y1 = 6 now just sub your info into the formula....do not try to solve it, and leave the parenthesis in there.

OpenStudy (anonymous):

so it 3x+y=3 @texaschic101

OpenStudy (texaschic101):

no.... y - y1 = m(x - x1) <-- this is the form you need y - 6 = -3(x - (-1) y - 6 = -3(x + 1) <-- your answer